In order to estimate the mean amount of time computer users spend on the internet each​ month, how many computer users must be surveyed in order to be 90​% confident that your sample mean is within 11 minutes of the population​ mean? Assume that the standard deviation of the population of monthly time spent on the internet is 208 min. What is a major obstacle to getting a good estimate of the population​ mean? Use technology to find the estimated minimum required sample size.

The minimum sample size required is [nothing] computer users.
Answer #1

solution Here population s.D() = 208 -12 1.6us 1968 error (e) = 11 At qoy confidence Zal = 1.645 sample size (n) = (20/2 = ho